Grade 11General Physics

The ratio of orbital velocity to escape velocity is

Hussayn Khan
8 Years agoGrade 11

1 Answer

Arun
8 Years ago
Dear student
 
Orbital speed at the surface is given by 

GM/R^2 = V^2 / R 

V = √ G M / R 

For escape velocity 

1/2 . m . V(e)^2 = GMm / R 

V(e) = √ 2 . G . M / R 

V(e) / V = √2 

( This is the same for any satellite around any planet or star )
 
 
 
Regards
Arun (askIITians forum expert)